R10 3300 R5 4700 R9 2200 R3 -V1 36800 -20 V ER2 310000 Vx R6 8200
Calculate voltages VR2, VR3, VRS, VR6, VR9, VR10, and Vx: Calculate currents 12, 13, 15, 16, 19, and 110- -
Answer #1

Selution Rio=33on R5 Ra 4ton 220 vo 1000 Z R₂ 68on - R6 82on Let Rake Resistor Rg and Rio are in parallel Combination of thes470 SR 20v Ry and R2 are in barallel. Let Rz lee Combination of these resistors. Rzz (RyllR2) = (1632 ||100) n. Rz= 620.06 aIves. = 8.62v Applying current division rule in above circuit. In = Is x Ry - 0.01834 x 1632 - Rat Ry 1000 + 1632 ---I2 0·011VR6 = I6 R6 = 6.97810 3 x 820 [VR6 = 5.7154 V By current division rule Ijo - Iyx Rq Rq+Rio Iso = 6.97X163 x 220 220+ 330 Ilo> + 20v VR2 R3 from the above circuit ka - VRzt VR6 Vx= 4.7396 +5.7154 Tx=10.455 V